the height of a triangle measures 9 inches more then its base, the area is 45 square inches, the measures of the base and height of the triangle?how do i find
Answer by checkley79(3341)

AREA=BH/2
H=B+9
45=B(B+9)/2
45=(B^2+9B)/2 CROSS MULTIPLY.
45*2=B^2+9B
90=B^2+9B
B^2+9B-90=0
(B+15)(B-6)=0
B-6=0
B=6 ANS. FOR THE BASE.
H=6+9=15 ANS. FOR THE HEIGHT.
PROOF:
45=6*15/2
45=90/2
45=45